OpenStudy (anonymous):

The total volume of two spheres is 10π cubic units. The ratio of the areas is 4:9. What is the volume of the smaller sphere in cubic units

OpenStudy (anonymous):

if you compare to the radius, the area goes with the square and the volume goes with the cube

OpenStudy (anonymous):

\[\frac{\text{small area}}{\text{large area}}=\frac{4}{9}\] and therefore \[\frac{\text{small radius}}{\text{large radius}}=\frac{2}{3}\]

OpenStudy (anonymous):

finally \[\frac{\text{small volume}}{\text{large volume}}=\frac{2^3}{3^3}\]